A particle travels so that its acceleration is given by a with rightwards harpoon with barb upwards on top equals 2 e to the power of negative 2 t end exponent space i with hat on top minus 3 cos t space j with hat on top space plus sin 2 t space k with hat on top




1
Expert's answer
2022-04-18T17:31:34-0400

"\vec a=2e^{-2t}\vec i-3\cos t\vec j+\sin 2t\vec k,"

"\vec v=-e^{-2t}\vec i-3\sin t\vec j-0 .5\cos 2t\vec k."